Which equation shows the substitution method being used to solve the system of linear equations?

x + y = 6
x = y + 5


A.
x = (x – 6) + 5

B.
x + (y + 5) = 6

C.
(y + 5) + y = 6

D.
x + y = y + 5

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

x + y = 6
x = y + 5
---
by substitution:
x + y = 6
(y + 5) + y = 6
---
answer:
C. (y + 5) + y = 6
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by stanbon(75887)

Which equation shows the substitution method being used to solve the system of linear equations?
x + y = 6
x = y + 5
-----------------------
Answer:: C.
(y + 5) + y = 6
=================
Cheers,
Stan H.
---------------